# DiPT: Enhancing LLM Reasoning through Diversified Perspective-Taking

### v1

@@ 4.3 Impact of DiPT-Enriched Fine-Tuning
We consider four models for training: Mistral7B-v0.1, Mistral7B-Instruct-v0.2, Llama3-8B, and Llama3-8B-Instruct. These models are fine-tuned on four distinct datasets, each representing a different task domain: OpenbookQA (common knowledge and understanding), GSM8K (grade school math word problems), CoQA (conversational dataset), and XWinograd (multilingual schema challenges). We evaluate the models’ performance in two settings, to assess their in-distribution and out-of-distribution generalization capabilities. For the in-distribution evaluation, we use the respective test split of the training distribution it was trained on to calculate the model’s performance. For the out-of-distribution (in-domain) evaluation, we use other datasets from a similar task domain to evaluate the model’s performance on data outside the training distribution but within the same domain. We group datasets into following domains: language understanding and knowledge (OpenbookQA, MMLU, PIQA, Hellaswag, Lambada), mathematical reasoning (GSM8K, MultiArith, SVAMP, AddSub), commonsense reasoning (CoQA, WSC, Winogrande, ARC-challenge), and multilingualism (XWinograd, WMT16, Lambada-multilingual). We train each model with the original dataset (plain), the CoT version of the dataset, or the DiPT+ CoT version of the dataset, using $3{,}000$ samples for each experiment, except for the base model. For further experimental details and all metrics, please refer to Appendix A.
@@ A.1 Tasks for Inference Stage
Zhang et al. [2015]. The AG News dataset is a collection of news articles categorically labeled into four classes (World, Sports, Business, and Science/Technology), providing a resource for text classification and topic modeling tasks. As news can belong to more than one category, we use top-2 accuracy.
@@
Socher et al. [2013]. The SST-5 dataset is a sentiment analysis dataset consisting of movie reviews categorized into five sentiment classes, including very negative, negative, neutral, positive, and very positive. We use a top-2 accuracy across methods because a sentiment might lie between 2 neighboring classes due to interpretation.
@@
Auer et al. [2007]. The DBpedia dataset is a knowledge base extracted from Wikipedia, representing structured information about a wide range of entities, including persons, places, organizations, and abstract concepts. We use top-1 accuracy.
@@
Huang et al. [2019]. The CosmosQA dataset is a reading comprehension dataset requiring contextual commonsense reasoning. The questions are posed as multi-choice problems that ask about likely causes or effects of events. We use top-1 accuracy.
@@
Li and Roth [2002], Hovy et al. [2001]. The TREC dataset is a question type classification dataset, which contains 6 coarse class labels. We use top-2 accuracy as the question type might belong to more than one category.
@@
Patel et al. [2021a]. The SVAMP dataset is consists of elementary-level math word problems. The dataset consists variations of the problems to test the model’s sensitivity to question understanding. Since the provided dataset is a single-answer dataset, we created three neighboring answers in addition to the groundtruth answer to make the problems multi-choice. We use top-1 accuracy.
@@
Lin et al. [2022]. The TruthfulQA dataset is used to measure the truthfulness of the model’s output generation. These problems are prone to be incorrectly answered if fallen into wrong beliefs and require correct pretrained information to be answered. We use top-1 accuracy.
@@
Cooper et al. [1996], Dagan et al. [2005]. The RTE dataset tests the language model in recognizing textual entailment in the provided context. The classification is binary. We use top-1 accuracy.
@@ A.2 Tasks for Fine-Tuning Stage
For these tasks, we use the popular evaluation repository LM Evaluation Harness to evaluate results for the following tasks [Gao et al., 2023].
@@
  • OpenbookQA [Mihaylov et al., 2018] - dataset designed to evaluate a model’s ability to apply elementary science knowledge to answer questions. We use the normalized top-1 accuracy.
  • MMLU [Mihaylov et al., 2018] - a comprehensive dataset encompassing a wide range of subjects to assess a model’s understanding across various academic disciplines and professional domains. We use the top-1 accuracy.
  • PIQA [Bisk et al., 2020] - a dataset that tests a model’s commonsense knowledge about the physical world. We use the normalized top-1 accuracy.
  • Hellaswag [Zellers et al., 2019] - a challenging dataset for commonsense reasoning, focusing on completing sentences in a way that makes sense in context. We use the normalized top-1 accuracy.
  • LAMBADA [Paperno et al., 2016] - a dataset designed to evaluate the ability of language models to understand and predict a missing word in a passage. We use the top-1 accuracy.
@@
  • GSM8K [Cobbe et al., 2021] - a dataset containing $8{,}000$ high-quality grade school math word problems designed to test arithmetic reasoning. We use the normalized 0-shot exact matching (flexible) accuracy.
  • MultiArith [Roy and Roth, 2016] - a dataset focused on arithmetic word problems that require multiple steps to solve. We use the normalized 0-shot exact matching (flexible) accuracy.
  • SVAMP (Single Variable Arithmetic Multiple Problems) [Patel et al., 2021b] - a dataset created to assess the robustness of models on arithmetic word problems. We use the normalized 0-shot exact matching (flexible) accuracy.
  • AddSub [Mishra et al., 2022] - a dataset consisting of arithmetic word problems that involve simple addition and subtraction. We use the normalized 0-shot exact matching (flexible) accuracy.
@@
  • CoQA (Conversational Question Answering) [Reddy et al., 2019] - dataset is designed for building conversational question answering systems. We use the F1 score.
  • WSC (The Winograd Schema Challenge)  [Levesque et al., 2012] - a dataset testing commonsense reasoning by identifying pronouns. We use the top-1 accuracy.
  • Winogrande [ai2, 2019] - a dataset extending WSC with more diverse and challenging sentences. We use the top-1 accuracy.
  • ARC Challenge (AI2 Reasoning Challenge) [Clark et al., 2018] - a dataset comprising of difficult multiple-choice science questions. We use the normalized top-1 accuracy.
@@
  • XWinograd [Muennighoff et al., 2022, Tikhonov and Ryabinin, 2021] - a multilingual version of the Winograd Schema Challenge. We use the top-1 accuracy.
  • WMT16 [Bojar et al., 2016] - a dataset consisting of parallel corpora and evaluation data for machine translation tasks. We report CHRF, BLEU, and TER scores and we use the CHRF [Popović, 2015] score accuracy for calculating the domain performance.
  • LAMBADA Multilingual [Paperno et al., 2016] - a dataset extending the original LAMBADA dataset to multiple languages. We use the top-1 accuracy.
@@ A.3 Hyperparameters
In our fine-tuning experiments, we train four models Mistral7B-v0.1, Mistral-7B-Instruct-v0.2 [Jiang et al., 2023], Llama3-8B, and Llama3-8B-Instruct [AI@Meta, 2024] with three Nvidia A100 80G GPUs. We follow the hyperparameter setup from Ethayarajh et al. [2023]. As such, we use a batch size of 32 and train for a single epoch. We keep the learning rate to be $75e-$ 7 as implemented. The maximum sequence length is set to 2048. We use RMSprop as our optimizer with warmup stages for 150 steps. The mixed precision is bfloat16.
@@ B.3 Results on Mistral 7B
We demonstrate the results on the open-weight model mistral-7b-instruct-v0.1 available on the HuggingFace library, which is a fine-tuned version on instructions of the mistral-7b-v0.1 model Jiang et al. [2023]. We present results in Table 10.
@@
0-Shot K-Shot ICL 0-Shot CoT Rephrasing Analogical Reasoner Ours

### v2

@@ 4.3 Impact of DiPT-Enriched Fine-Tuning
We consider four models for training: Mistral7B-v0.1, Mistral7B-Instruct-v0.2, Llama3-8B, and Llama3-8B-Instruct. These models are fine-tuned on four distinct datasets, each representing a different task domain: OpenbookQA (common knowledge and understanding), GSM8K (grade school math word problems), and CoQA (conversational dataset). We evaluate the models’ performance in two settings, to assess their in-distribution and out-of-distribution generalization capabilities. For the in-distribution evaluation, we use the respective test split of the training distribution it was trained on to calculate the model’s performance. For the out-of-distribution (in-domain) evaluation, we use other datasets from a similar task domain to evaluate the model’s performance on data outside the training distribution but within the same domain. We group datasets into following domains: language understanding and knowledge (OpenbookQA, MMLU, PIQA, Hellaswag, Lambada), mathematical reasoning (GSM8K, MultiArith, SVAMP, AddSub), and commonsense reasoning (CoQA, WSC, Winogrande, ARC-challenge). We train each model with the original dataset (plain), the CoT version of the dataset, or the DiPT+ CoT version of the dataset, using $3{,}000$ samples for each experiment, without mixing data between different data types to ensure fair comparison. For further experimental details and all metrics, please refer to Appendix A.
@@ 7 Ethical Considerations
8 Acknowledgments
Ruoxi Jia and the ReDS lab acknowledge support through grants from the Amazon-Virginia Tech Initiative for Efficient and Robust Machine Learning, the Cisco Award, the Commonwealth Cyber Initiative Cybersecurity Research Award, the National Science Foundation under grants IIS-2312794, IIS-2313130, OAC-2239622, and OpenAI API research credits.
@@ A.1 Tasks for Inference Stage
Zhang et al. (2015). The AG News dataset is a collection of news articles categorically labeled into four classes (World, Sports, Business, and Science/Technology), providing a resource for text classification and topic modeling tasks. As news can belong to more than one category, we use top-2 accuracy.
@@
Socher et al. (2013). The SST-5 dataset is a sentiment analysis dataset consisting of movie reviews categorized into five sentiment classes, including very negative, negative, neutral, positive, and very positive. We use a top-2 accuracy across methods because a sentiment might lie between 2 neighboring classes due to interpretation.
@@
Auer et al. (2007). The DBpedia dataset is a knowledge base extracted from Wikipedia, representing structured information about a wide range of entities, including persons, places, organizations, and abstract concepts. We use top-1 accuracy.
@@
Huang et al. (2019). The CosmosQA dataset is a reading comprehension dataset requiring contextual commonsense reasoning. The questions are posed as multi-choice problems that ask about likely causes or effects of events. We use top-1 accuracy.
@@
Li and Roth (2002); Hovy et al. (2001). The TREC dataset is a question type classification dataset, which contains 6 coarse class labels. We use top-2 accuracy as the question type might belong to more than one category.
@@
Patel et al. (2021a). The SVAMP dataset is consists of elementary-level math word problems. The dataset consists variations of the problems to test the model’s sensitivity to question understanding. Since the provided dataset is a single-answer dataset, we created three neighboring answers in addition to the groundtruth answer to make the problems multi-choice. We use top-1 accuracy.
@@
Lin et al. (2022). The TruthfulQA dataset is used to measure the truthfulness of the model’s output generation. These problems are prone to be incorrectly answered if fallen into wrong beliefs and require correct pretrained information to be answered. We use top-1 accuracy.
@@
Cooper et al. (1996); Dagan et al. (2005). The RTE dataset tests the language model in recognizing textual entailment in the provided context. The classification is binary. We use top-1 accuracy.
DROP (Discrete Reasoning Over Paragraphs)
Dua et al. (2019). The DROP dataset is a benchmark designed to test reading comprehension by requiring discrete reasoning, such as numerical operations and logical inferences, over diverse paragraphs of text. We use exact matching.
MATH
Hendrycks et al. (2021). The MATH dataset is a benchmark dataset designed to evaluate the mathematical reasoning and problem-solving capabilities of AI models, containing high school-level math problems across various domains such as algebra, calculus, and geometry. We sample Level 4 and 5 difficulty problems. We use exact matching.
GPQA (Google-Proof QA)
Rein et al. (2023). GPQA benchmark is a challenging question-answering benchmark dataset designed to test AI models on graduate-level topics across various academic disciplines, with questions that are difficult to answer through simple web searches. We use exact matching.
@@ A.2 Tasks for Fine-Tuning Stage
For these tasks, we use the popular evaluation repository LM Evaluation Harness to evaluate results for the following tasks (Gao et al., 2023).
@@
  • OpenbookQA (Mihaylov et al., 2018) - dataset designed to evaluate a model’s ability to apply elementary science knowledge to answer questions. We use the normalized top-1 accuracy.
  • MMLU (Mihaylov et al., 2018) - a comprehensive dataset encompassing a wide range of subjects to assess a model’s understanding across various academic disciplines and professional domains. We use the top-1 accuracy.
  • PIQA (Bisk et al., 2020) - a dataset that tests a model’s commonsense knowledge about the physical world. We use the normalized top-1 accuracy.
  • Hellaswag (Zellers et al., 2019) - a challenging dataset for commonsense reasoning, focusing on completing sentences in a way that makes sense in context. We use the normalized top-1 accuracy.
  • LAMBADA (Paperno et al., 2016) - a dataset designed to evaluate the ability of language models to understand and predict a missing word in a passage. We use the top-1 accuracy.
@@
  • GSM8K (Cobbe et al., 2021) - a dataset containing $8{,}000$ high-quality grade school math word problems designed to test arithmetic reasoning. We use the normalized 0-shot exact matching (flexible) accuracy.
  • MultiArith (Roy and Roth, 2016) - a dataset focused on arithmetic word problems that require multiple steps to solve. We use the normalized 0-shot exact matching (flexible) accuracy.
  • SVAMP (Single Variable Arithmetic Multiple Problems) (Patel et al., 2021b) - a dataset created to assess the robustness of models on arithmetic word problems. We use the normalized 0-shot exact matching (flexible) accuracy.
  • AddSub (Mishra et al., 2022) - a dataset consisting of arithmetic word problems that involve simple addition and subtraction. We use the normalized 0-shot exact matching (flexible) accuracy.
@@
  • CoQA (Conversational Question Answering) (Reddy et al., 2019) - dataset is designed for building conversational question answering systems. We use the F1 score.
  • WSC (The Winograd Schema Challenge)  (Levesque et al., 2012) - a dataset testing commonsense reasoning by identifying pronouns. We use the top-1 accuracy.
  • Winogrande (ai2, 2019) - a dataset extending WSC with more diverse and challenging sentences. We use the top-1 accuracy.
  • ARC Challenge (AI2 Reasoning Challenge) (Clark et al., 2018) - a dataset comprising of difficult multiple-choice science questions. We use the normalized top-1 accuracy.
@@
  • XWinograd (Muennighoff et al., 2022; Tikhonov and Ryabinin, 2021) - a multilingual version of the Winograd Schema Challenge. We use the top-1 accuracy.
  • WMT16 (Bojar et al., 2016) - a dataset consisting of parallel corpora and evaluation data for machine translation tasks. We report CHRF, BLEU, and TER scores and we use the CHRF (Popović, 2015) score accuracy for calculating the domain performance.
  • LAMBADA Multilingual (Paperno et al., 2016) - a dataset extending the original LAMBADA dataset to multiple languages. We use the top-1 accuracy.
@@ A.3 Hyperparameters
In our fine-tuning experiments, we train four models Mistral7B-v0.1, Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Llama3-8B, and Llama3-8B-Instruct (AI@Meta, 2024) with three Nvidia A100 80G GPUs. We follow the hyperparameter setup from Ethayarajh et al. (2023). As such, we use a batch size of 32 and train for a single epoch. We keep the learning rate to be $75e-$ 7 as implemented. The maximum sequence length is set to 2048. We use RMSprop as our optimizer with warmup stages for 150 steps. The mixed precision is bfloat16.
We used the default hyperparameters for decoding for GPT-4, where the presence penalty is set to 0, the temperature to 1, top-p to 1, and the frequency penalty to 0.
@@ Appendix B Additional Results
Dataset Labeling Error Detection.
We observe that although our method achieves high performance, it cannot reach $100\%$. Upon closer examination, we identified errors within datasets that prevented our method from achieving a perfect score. Current works in reasoning often use datasets commonly employed in NLP. However, for a dataset to serve as a reliable benchmark, it must exhibit high quality without errors. Otherwise, achieving high performance on inaccurately labeled data can mislead the comprehension of the method. Consequently, we aim to prevent such errors in these datasets. We apply our method to verify the labeling of these datasets and identify potential errors in the misalignment of the labels. Specifically, we employ DiPT on the gpt-4-1106-preview model to identify mismatched labels between the predicted and annotated labels. Then, to evaluate the correctness of error identification by our method, we leverage expertise evaluations from several powerful LLMs, including Bard/Gemini and Claude, in conjunction with our judgments. Using Krippendorff’s Alpha to measure the Inter-Annotator Agreement between all raters (LLMs and humans), we reached an alpha of 0.67, indicating a strong agreement between raters, and 0.89 between humans alone. We use fine-grained metrics to better categorize the labeling errors: Wrong label, where all experts disagree with the original ground truth label, Ambiguous examples, where some experts disagree with the original label, and False Positives, where all experts agree with the original label.
SST-5
AG News
TREC
DBPedia
CosmosQA
SVAMP
TruthfulQA
RTE
Wrong 15 4 5 1 5 5 8 3
Ambiguous 4 0 2 0 6 0 3 5
False Positive 1 0 2 0 2 0 10 2
@@ B.3 Results on Mistral 7B
We demonstrate the results on the open-weight model mistral-7b-instruct-v0.1 available on the HuggingFace library, which is a fine-tuned version on instructions of the mistral-7b-v0.1 model Jiang et al. (2023). We present results in Table 13.
@@
0-Shot K-Shot ICL 0-Shot CoT Rephrasing Analogical Reasoner DiPT